solved the equation in standard form
x^2 - 3x - 4 = 0

Answer by ewatrrr(24785)


Hi,
Solving
x^2 - 3x - 4 = 0 OR y = (x-3/2)^2 - 25/4 (vertex form by completing square)
factoring
(x-4)(x+1)=0 Note:SUM of the inner product(-4x) and the outer product(1*x) = -3x
(x-4)=0 x = 4
(x+1)=0 x = -1